OpenStudy (anonymous):

what is 2 rad 15 in simplest foarm

jimthompson5910 (jim_thompson5910):

\[\Large 2\sqrt{15}\] is as simple as it gets

jimthompson5910 (jim_thompson5910):

you can't simplify that any further

OpenStudy (anonymous):

2 rad 45 is the question, i put the wrong number in

jimthompson5910 (jim_thompson5910):

which perfect squares are factors of 45

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

so 45 = 9*5

jimthompson5910 (jim_thompson5910):

which means \[\Large \sqrt{45} = \sqrt{9*5}\] \[\Large \sqrt{45} = \sqrt{9}*\sqrt{5}\] \[\Large \sqrt{45} = 3*\sqrt{5}\]

jimthompson5910 (jim_thompson5910):

which tells us that \[\Large 2\sqrt{45}\] turns into \[\Large 2*3\sqrt{5}\] \[\Large 6\sqrt{5}\] So in the end, \[\Large 2\sqrt{45} = 6\sqrt{5}\]

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np